The intensity of solar radiation reaching Mars averages about 580 W/m^2.

Assuming that Mars behaves like a blackbody, how much energy is absorbed per unit time. What is equilibrium temperature if all this energy is re-radiated back into space?
Answer #1

a)
radius of mars, r = 3.39*10^6 m

surface area of mars = 4*pi*r^2

= 4*pi*(3.39*10^6)^2

= 1.44*10^14 m^2

Enery obsorbed per seccond = Intensity*area

= 580*1.44*10^14

= 8.352*10^16 J <<<<<-------Answer

b)

Use stefan's law

Eergy emited by a body per unit area per unit time,
I = sigma*T^4

T = (I/sigma)^(1/4)

= (580/(5.67*10^-8))^(1/4)

= 318 k <<<<<-------Answer